# From Ground to Sky: Architectures, Applications, and Challenges Shaping Low-Altitude Wireless Networks

### v2

@@ Abstract
In this article, we introduce a novel low-altitude wireless network (LAWN), which is a reconfigurable, three-dimensional (3D) layered architecture. In particular, the LAWN integrates connectivity, sensing, control, and computing across aerial and terrestrial nodes that enable seamless operation in complex, dynamic, and mission-critical environments. Different from the conventional aerial communication systems, LAWN’s distinctive feature is its tight integration of functional planes in which multiple functionalities continually reshape themselves to operate safely and efficiently in the low-altitude sky. With the LAWN, we discuss several enabling technologies, such as integrated sensing and communication (ISAC), semantic communication, and fully-actuated control systems. Finally, we identify potential applications and key cross-layer challenges. This article offers a comprehensive roadmap for future research and development in the low-altitude airspace.
Index Terms:
Low-altitude wireless networks, data transmission, environmental sensing, low-latency control.
@@ I Introduction
The low-altitude airspace, roughly below $3,000$m above ground level, has rapidly evolved into a third dimension for digital infrastructure supporting logistics, agriculture, and public safety. It is projected that by 2030, the commercial drone sector will have a value of USD \$$\$260$ billion, four times higher than today’s state[1]. In parallel, regulators are clearing the skies: the U.S. Federal Aviation Administration (FAA)’s Beyond-Visual-Line-of-Sight rule-making committee issued final recommendations in 2024 that pave the way for routine long-range flights without direct pilot oversight, while Europe’s U-space framework harmonized cooperative traffic management at $120$m above ground level (AGL). At the same time, the civil aviation administration of China (CAAC) made advanced air-mobility and “low-altitude economy” a national plan, charting a phased roadmap for exploiting low-altitude airspace. The convergence of market growth, permissive regulation, and maturing autonomy is giving rise to the low-altitude wireless network (LAWN). By eliminating the constraint of fixed ground stations and enabling rapid topology re-configuration, LAWNs promise various capabilities for precision agriculture, distributed aerial sensing, swarm-based disaster response, and the broader low-altitude economy, as shown in Fig. 1.
@@
While aerial communications have been investigated extensively in the last decade, the LAWN is more than the concept of “cellular-in-the-sky” [2]. That is, the LAWN is a flexible, three-dimensional web that integrates connectivity, real-time sensing, and closed-loop control across multiple drones and terrestrial nodes. Unlike terrestrial or high-altitude systems, the operational environment for the LAWN is usually densely cluttered (e.g., $100$ aerial drones per km2) and rapidly evolving. For instance, multiple drones collaborate to execute specific tasks and may take various actions such as climbing, hovering, and banking at up to $30$m/s velocity. Moreover, the clear line-of-sight (LOS) drone-to-drone or drone-to-ground links may translate into non-line-of-sight (NLOS) behind buildings. Therefore, the signal paths, interference patterns, and safety envelopes may change in a very short time duration. Moreover, the flight control often demands end-to-end latency below $10$ms with $99.999$% reliability to achieve the stability[3]. To satisfy the combined needs of joint connectivity-sensing–control and even computing, a LAWN is expected to be elastic in both topology and function: the drones within range should be capable of sharing information without human control; the radio waveform carrying data may be used for radar sensing such that the drone can “see” obstacles while it talks; the on-board drone controllers could reroute or re-plan the swarm formation to combat environmental dynamics, and allocate computing resources between drones and edge servers for balancing battery consumptions. Essentially, a LAWN is a cross-disciplinary concept in which multiple functionalities continually reshape themselves to operate safely and efficiently in the low-altitude sky.
@@
Despite the great potential of LAWN for the 6G ecosystem, most prior works tackle isolated topics such as drone-to-ground communication and trajectory design. This article fills this gap by offering the first end-to-end perspective on what a LAWN is, why it matters, and how it can be realized. In particular, we commence from the architectures of a LAWN by presenting its functional planes, i.e., data transmission, control, sensing, and computing, clarifying where it diverges from traditional aerial communications or drone-assisted mesh networks. Next, we highlight several enabling technologies, such as integrated sensing and communication (ISAC) and artificial intelligence (AI) and potential applications, e.g., precision agriculture related to the LAWN. Then, a case study of swarm coordination in post-disaster scenarios is discussed. To promote the realization of the LAWN, we summarize various challenges, such as 3D spectrum management, cyber-physical security, and regulations. Finally, we draw the conclusions.
@@
Figure 2: The network architecture of a LAWN. There are three tightly coupled functional planes, i.e., the data plane, the control plane, and the sensing plane to provide the capabilities of data transmission, environmental awareness, and flight control. An auxiliary computing plane further supports data processing and decision-making.
@@ II LAWN Architectures: A 3-D Web of Connectivity, Sensing, and Control
The aerial drone-enabled communications can be traced back to mid 1990s, when drones were considered to serve as platforms for a high capacity trunk radio relay[2]. Although these “flying microwave towers” proved the value of airborne connectivity, their large sizes, high costs, and single-purpose payloads limited wider adoption. Since 2010, the rapid advances in lithium batteries, carbon-composite structures, and system-on-chip radios boosted the research on drone communications. The major benefit of drone communications lies in its flexibility for providing reliable LoS links. Therefore, most research to date regards drones purely as airbone base stations or relays to optimize spectral efficiency, mitigate ground-user interference, or reduce drone-to-ground path loss. Nevertheless, future low-altitude missions, such as autonomous delivery, cooperative cinematic swarms, and electric vertical take-off and landing (eVTOL) passenger traffic demands much more than data transmission. They demand real-time sensing, distributed decision-making, swarm coordination, and aviation-grade safety assurance. These multi-dimensional requirements motivate the concept of LAWN, which is to be discussed in the following.
@@ II-A What is the LAWN?
The LAWN refers to a dynamically reconfigurable three-dimensional network architecture that integrates aerial nodes operating at altitudes typically below $3,000$ m and terrestrial nodes to form a unified platform. This architecture allows for high-speed data transmission ($10-100$Mbps), ultra-reliable low-latency command-and-control ($<10^{-5}$ error rate and $10$ms latency), and high-precision sensing (centimetre-level accuracy). The LAWN serves multiple functions, including delivery and logistics for commercial applications, as well as search and rescue operations in disaster scenarios. Fig. 2 provides a visual overview of the LAWN architecture. At the heart of the LAWN are three tightly coupled functional planes (i.e., the Data Plane, the Control Plane, and the Sensing Plane), as well as an auxiliary Computing Plane.
@@ II-A1 Data Plane
The data plane aims to provide on-demand communications, which is responsible for transmitting the mission payloads, such as videos, mapping information, software updates, between aerial drones and terrestrial nodes in the LAWN[2]. To do so, the data plane supports three fundamental links, i.e.,
  • Drone-to-Ground (D2G) link: Mission payloads such as high‑quality video or images, as well as other required information must reach ground operators or edge servers in near real time. The D2G link therefore acts as the primary link of the data plane.
  • Ground-to-Drone (G2D) link: It carries uplink information to the drones, typically including navigation maps, software or model updates, and flight plans. Because the uplink information might arrive in bursts, the G2D link is provisioned as a scheduling priority.
  • Drone-to-Drone (D2D) link: It enables each drone to share the sensor data as well as the swarm coordination message to neighboring drones. The D2D connectivity transforms individual drones into a flexible and cooperative mesh, extending coverage beyond any single drone and ensuring robustness when executing tasks.
Although LAWN primarily operates within the low-altitude airspace, high-altitude platforms (HAPs) and low-earth orbit (LEO) satellites can serve as supplementary communication nodes. They provide extended backhaul or remote relay capabilities, especially in remote regions lacking terrestrial infrastructure.
Due to the realities of low‑altitude airspace, there are some common impairments across the data plane of LAWN. A drone can yaw tens of degrees per second and cruise at 1030103010–3010 – 30m/s, turning a clear LoS link to a NLoS one. A data plane scheduler is therefore needed to continuously tag each packet burst with LoS/NLoS status, instantaneous path‑loss, and blockage probability [4]. If the tag indicates a positive path state, the data keeps flowing; if a building blocks the signal with a high chance, the scheduler may choose to divert the packets to a nearby relay drone or only transmit the small and time‑critical packets. The relative velocities between different nodes introduce the Doppler shift up to kilohertz, depending on the carrier frequency. In such cases, the data plane will allow coding, hybrid automatic repeat request, and packet‑aggregation timers to shrink on fast‑fading channels. Moreover, for the LAWN operating at heterogeneous bands, e.g., mixed sub-6 GHz and mmWave bands, the data plane could dynamically resize channel bands and link transmit powers by performing the make‑before‑break handoffs.
@@ II-A2 Control Plane
The control plane can be seen as the nervous system of the LAWN, aiming for keeping the sky orderly and safe, and therefore having the highest priority for radio resource sharing[3]. On the one hand, it guarantees that every drone in the sky is uniquely identified and properly separated, ensuring the safety to fly. On the other hand, the control plane also enables multiple drones to maneuver in a coordinated group manner. Meeting these goals demands stricter service requirements than the data plane, e.g., a latency of $<10$ms and reliability of $>99.999$%, as well as cryptographic authenticity for each packet.
@@ II-A3 Sensing Plane
  • Radio Frequency (RF) Sensing: In addition to passive sensors, the LAWN nodes can exploit radar-like capabilities using radio signals to estimate range and Doppler shift of surrounding objects. Radar waveforms measure range and Doppler shifts to determine the position and velocity of objects or other nodes, performing effectively in both LoS and NLoS conditions. When integrated with D2D and D2G links, RF sensing enables cooperative sensing by sharing data across the network.
@@
II-A4 Auxiliary Computing Plane
In addition to the three primary functional planes above, the computing plane serves as an auxiliary functional layer within LAWNs, which is dedicated to providing the computational resources required for real-time data processing, signal analysis, and decision-making. Its auxiliary role enhances the primary functions of connectivity, control, and sensing in LAWNs. The computing plane is composed of on-board processors, edge servers, and cloud-based infrastructure.
At the node level, on-board processors enable immediate execution of time-sensitive tasks such as obstacle detection, and self path planning. For more computationally intensive tasks, such as map reconstruction, the computing plane could leverage edge offloading, dynamically assigning workloads from resource-constrained drones to nearby edge nodes[6]. The edge servers, typically deployed at ground base stations, provide substantial computational resources for data processing. Through dynamic task offloading, the computing plane can reduce onboard computational load while meeting latency requirements. Moreover, cloud-based infrastructure provides high-capacity computing and long-term storage. While cloud nodes are not suitable for latency-critical operations, they are essential for performing analysis and network optimization. By integrating onboard, edge, and cloud resources, the computing plane of LAWNs supports a distributed computing framework that enables flexible task allocation across the network, which enhances network scalability and resilience, particularly in large-scale deployments.
We can see that all four planes are closely interconnected, working together to support the LAWN’s adaptive capabilities. For instance, sensing data about a nearby building informs the control plane to reroute flight paths instantly. Meanwhile, the data plane ensures that this sensing data, along with C2 signals, is transmitted efficiently, while maintaining network-wide awareness. The computing plane will process raw sensing data into actionable insights, optimize control algorithms, and balance computational loads between drones and edge servers. This tight integration allows the LAWN to respond seamlessly to rapidly changing conditions, supporting applications such as swarm coordination, disaster response, urban surveillance, package delivery, and environmental monitoring.
@@ II-B How LAWN Differs from Aerial Communications?
Functional Scope Focused on communication only; separate control and sensing layers Integrated data, control, and sensing planes with a supporting computing layer
@@
Computing Resources Centralized or onboard-only processing; limited scalability Hierarchical computing (onboard, edge, and cloud) for real-time processing and decision-making
@@
In contrast, the LAWN introduces a paradigm shift by redefining the network around the terrestrial infrastructure, extending its functionality into the low-altitude space through a reconfigurable, intelligent, and mission-aware network “skyway”. Instead of simply using drones as communication tools, LAWN treats aerial and terrestrial nodes as jointly organized components of a cyber-physical system. The architecture tightly integrates three primary functional planes, i.e., data, control, and sensing, augmented by an auxiliary computing plane. For example, traditional systems relying on fixed ground stations may struggle to adapt to such dynamic conditions. Dynamic reconfigurability of both aerial and terrestrial nodes sets LAWN apart from the configurations of traditional systems, ensuring robustness across diverse scenarios. A defining advantage of LAWN is its cooperative sensing capability, allowing aerial and ground nodes to share environmental data in real time. This shared perception supports safety-critical services such as multi-drone formation flying, dynamic obstacle avoidance, and urban air traffic management [^1] Traditional drone-assisted mesh networks, while capable of network-wide sharing, often lack sensing functionality supported by the sensing plane of the LAWN.. Additionally, LAWN’s auxiliary computing plane provides distributed computational resources for real-time processing and decision-making. This includes on-board processing for immediate tasks, edge computing for latency-sensitive operations, and cloud-based infrastructure for intensive computations and long-term data analysis. Such a hierarchical computing framework is rarely embedded in traditional systems. In Table I, we summarize the comparison between the LAWN and the aerial communication system.

### v3

@@ Abstract
In this article, we introduce a novel low-altitude wireless network (LAWN), which is a reconfigurable, three-dimensional (3D) layered architecture. In particular, the LAWN integrates connectivity, sensing, control, and intelligence across aerial and terrestrial nodes that enable seamless operation in complex, dynamic, and mission-critical environments. Different from the conventional aerial communication systems, LAWN’s distinctive feature is its tight integration of functional planes in which multiple functionalities continually reshape themselves to operate safely and efficiently in the low-altitude sky. With the LAWN, we discuss several enabling technologies, such as integrated sensing and communication (ISAC), semantic communication, and fully-actuated control systems. Finally, we identify potential applications and key cross-layer challenges. This article offers a comprehensive roadmap for future research and development in the low-altitude airspace.
@@ I Introduction
The low-altitude airspace, roughly below 3,0003,000m above ground level, has rapidly evolved into a third dimension for digital infrastructure supporting logistics, agriculture, and public safety. It is projected that by 2030, the commercial drone sector will have a value of USD $\mathdollar 260$ billion, four times higher than today’s state[10]. In parallel, regulators are clearing the skies: the U.S. Federal Aviation Administration (FAA)’s Beyond-Visual-Line-of-Sight rule-making committee issued final recommendations in 2024 that pave the way for routine long-range flights without direct pilot oversight, while Europe’s U-space framework harmonized cooperative traffic management at 120120m above ground level (AGL). At the same time, the civil aviation administration of China (CAAC) made advanced air-mobility and “low-altitude economy” a national plan, charting a phased roadmap for exploiting low-altitude airspace. The convergence of market growth, permissive regulation, and maturing autonomy is giving rise to the low-altitude wireless network (LAWN). By eliminating the constraint of fixed ground stations and enabling rapid topology re-configuration, LAWNs promise various capabilities for precision agriculture, distributed aerial sensing, swarm-based disaster response, and the broader low-altitude economy, as shown in Fig. 1.
@@
While aerial communications have been investigated extensively in the last decade, a LAWN is more than the concept of “cellular-in-the-sky” [26]. A LAWN is a flexible three-dimensional web that integrates connectivity, real-time sensing, and closed-loop control across multiple drones and terrestrial nodes. Unlike terrestrial or high-altitude systems, the operational environment of a LAWN is relatively densely cluttered (e.g., up to $100$ aerial drones per km2) and rapidly evolving. For instance, drones may execute aggressive maneuvers such as climbing, hovering, and banking at velocities up to 30 m/s, while line-of-sight (LoS) drone-to-drone or drone-to-ground links can quickly turn into non-line-of-sight (NLoS) behind buildings. As a result, signal paths, interference patterns, and safety envelopes may change within a short duration. Meanwhile, flight control often demands end-to-end latency below $10$ ms with 99.999% reliability to ensure stability. Meeting these coupled requirements requires a LAWN to be elastic in both topology and function. Different drones within range should autonomously share information. The radio waveform carrying data may also be exploited for sensing so that a drone can “see” obstacles while it communicates. The controllers should reroute or re-plan swarm formations to combat environmental dynamics, while coordinating intelligence & computing resources across drones and edge servers to balance performance and battery consumption. In this sense, a LAWN is inherently cross-disciplinary, where communication, sensing, control, and intelligence continually reshape each other to sustain safe and efficient operation in the low-altitude sky.
@@
Despite the great potential of LAWN for the 6G ecosystem, most prior works address isolated topics such as drone-to-ground communication or trajectory design. This article fills this gap by offering a holistic tutorial perspective on what a LAWN is, why it matters, and how it can be realized. In particular, we commence from LAWN architectures by presenting its functional planes, namely data transmission, control, sensing, and intelligence&computing, clarifying where it diverges from traditional aerial communications or drone-assisted mesh networks. Next, we highlight enabling technologies such as integrated sensing and communication (ISAC) and artificial intelligence (AI), together with representative applications, e.g., precision agriculture. We then discuss a case study of swarm coordination in post-disaster scenarios. To promote the realization of LAWNs, we summarize key challenges including 3D spectrum management, cyber-physical security, and regulation. Finally, we draw conclusions and outline research directions. We hope that this paper will serve as a tutorial for researchers in this area and inspire low-altitude innovations.
@@
Figure 2: The network architecture of a LAWN. There are three tightly coupled functional planes, i.e., the data plane, the control plane, and the sensing plane to provide the capabilities of data transmission, environmental awareness, and flight control. An auxiliary intelligence&computing plane further supports distributed inference and hierarchical onboard–edge–cloud processing to enable timely decision-making.
@@ II LAWN Architectures: A 3-D Web of Connectivity, Sensing, and Control
The aerial drone-enabled communications can be traced back to mid 1990s, when drones were considered to serve as platforms for a high capacity trunk radio relay[26]. Although these “flying microwave towers” proved the value of airborne connectivity, their large sizes, high costs, and single-purpose payloads limited wider adoption. Since 2010, the rapid advances in lithium batteries, carbon-composite structures, and system-on-chip radios boosted the research on drone communications. The major benefit of drone communications lies in its flexibility for providing reliable LoS links. Therefore, most research to date regards drones purely as airborne base stations or relays to optimize spectral efficiency, mitigate ground-user interference, or reduce drone-to-ground path loss. Nevertheless, future low-altitude missions, such as autonomous delivery, cooperative cinematic swarms, and electric vertical take-off and landing (eVTOL) passenger traffic demands much more than data transmission. They demand real-time sensing, distributed decision-making, swarm coordination, and aviation-grade safety assurance. These multi-dimensional requirements motivate the concept of LAWN, which is to be discussed in the following.
@@ II-A What is the LAWN?
The LAWN refers to a dynamically reconfigurable three-dimensional network architecture that integrates aerial nodes operating at altitudes typically below $3,000$ m and terrestrial nodes to form a unified platform. This architecture allows for high-speed data transmission (1010010-100Mbps), ultra-reliable low-latency command-and-control ($10^{-5}$ error rate and 1010ms latency), and high-precision sensing (centimeter-level accuracy). The LAWN serves multiple functions, including delivery and logistics for commercial applications, as well as search and rescue operations in disaster scenarios. Fig. 2 provides a visual overview of the LAWN architecture. At the heart of the LAWN are three tightly coupled functional planes, i.e., the Data Plane, the Control Plane, and the Sensing Plane, as well as an auxiliary Intelligence&Computing (I&C) Plane.
@@ II-A1 Data Plane
The data plane aims to provide on-demand communications, which is responsible for transmitting the mission payloads, such as videos, mapping information, and software updates between aerial drones and terrestrial nodes in the LAWN[26]. To do so, the data plane supports three fundamental links, i.e.,
  • Air-to-Groud (A2G) link: Mission payloads such as high‑quality video or images, as well as other required information must reach ground operators or edge servers in near real time. The A2G link therefore acts as the primary link of the data plane.
  • Ground-to-Air (G2A) link: It carries uplink information to the drones, typically including navigation maps, software or model updates, and flight plans. Because the uplink information might arrive in bursts, the G2A link is provisioned as a scheduling priority.
  • Air-to-Air (A2A) link: It enables each drone to share the sensor data as well as the swarm coordination message to neighboring drones. The A2A connectivity transforms individual drones into a flexible and cooperative mesh, extending coverage beyond any single drone and ensuring robustness when executing tasks.
Operationally, the data plane decides who connects to whom, on which band, and with what redundancy, while adapting scheduling and retransmission to LoS/NLoS transitions and kHz Doppler. Although LAWN primarily operates within the low-altitude airspace, high-altitude platforms (HAPs) and low-earth orbit (LEO) satellites can serve as supplementary communication nodes. They provide extended backhaul or remote relay capabilities, especially in remote regions lacking terrestrial infrastructure.
Due to the realities of low‑altitude airspace, there are some common impairments across the data plane of LAWN. A drone can yaw tens of degrees per second and cruise at 103010–30m/s, turning a clear LoS link to a NLoS one. A data plane scheduler is therefore needed to continuously tag each packet burst with LoS/NLoS status, instantaneous path‑loss, and blockage probability [7]. If the tag indicates a positive path state, the data keeps flowing; if a building blocks the signal with a high chance, the scheduler may choose to divert the packets to a nearby relay drone or only transmit the small and time‑critical packets. The relative velocities between different nodes introduce the Doppler shift up to kilohertz, depending on the carrier frequency. In such cases, the data plane will allow coding, hybrid automatic repeat request, and packet‑aggregation timers to shrink on fast‑fading channels. Moreover, for the LAWN operating at heterogeneous bands, e.g., mixed sub-6 GHz and mmWave bands, the data plane could dynamically resize channel bands and link transmit powers by performing the make‑before‑break handoffs.
@@ II-A2 Control Plane
The control plane can be seen as the nervous system of the LAWN, enforcing aviation-grade safety including identity, separation, and contingency actions while enabling coordinated multi-drone maneuvering under stringent latency/reliability and authenticity constraints.[20]. On the one hand, it guarantees that every drone in the sky is uniquely identified and properly separated, ensuring the safety to fly. On the other hand, the control plane also enables multiple drones to maneuver in a coordinated group manner. Meeting these goals demands stricter service requirements than the data plane, e.g., a latency of <10<10ms and reliability of >99.999>99.999%, as well as cryptographic authenticity for each packet.
@@ II-A3 Sensing Plane
  • Radio Frequency (RF) Sensing: In addition to passive sensors, the LAWN nodes can exploit radar-like capabilities using radio signals to estimate range and Doppler shift of surrounding objects. A commonly used model is
    $r(t)=\sum_{p}\alpha_{p}\,x(t-\tau_{p})e^{j2\pi\nu_{p}t}+w(t),$ (1)
    where $x(t)$ is an known probing waveform and $(\tau_{p},\nu_{p})$ correspond to range and radial velocity of the $p$-th target, respectively. RF waveforms measure $\tau_{p}$ and $\nu_{p}$ to determine the position and velocity of objects or other nodes, performing effectively in both LoS and NLoS conditions. These estimates complement cameras/LiDAR/IMU and become more powerful when shared over A2A/A2G links for cooperative sensing.
@@
II-A4 Intelligence & Computing Plane
In addition to the three primary functional planes above, the I&C plane serves as an supporting functional layer within LAWNs, which is dedicated to providing the computational resources and distributed intelligence required for real-time data processing, signal analysis, and decision-making. Its auxiliary role enhances the primary functions of connectivity, control, and sensing in LAWNs. The I&C plane is composed of on-board processors, edge servers, and cloud-based infrastructure. Beyond providing computing resources, this plane also hosts distributed inference, prediction, and decision support that directly shapes scheduling and control.
At the node level, on-board processors enable immediate execution of time-sensitive tasks such as obstacle detection, and self path planning. For more computationally intensive tasks, such as map reconstruction, the I&C plane could leverage edge offloading, dynamically assigning workloads from resource-constrained drones to nearby edge nodes[11]. The edge servers, typically deployed at ground base stations, provide substantial resources for data processing. Through dynamic task offloading, the I&C plane can reduce onboard computational load while meeting latency requirements. Moreover, cloud-based infrastructure provides high-capacity computing, long-term storage, and model training. While cloud nodes are not suitable for latency-critical operations, they are essential for performing analysis and network optimization. By integrating onboard, edge, and cloud resources, the I&C plane of LAWNs supports flexible placement of both computation and intelligence across the network, improving scalability and resilience in large-scale deployments.
We can see that all four planes are closely interconnected, working together to support the LAWN’s adaptive capabilities. For instance, sensing data about a nearby building informs the control plane to reroute flight paths instantly. Meanwhile, the data plane ensures that this sensing data, along with C2 signals, is transmitted efficiently, while maintaining network-wide awareness. The I&C plane will process raw sensing data into actionable insights, supports distributed inference/learning, optimizes control algorithms, and balances workloads between drones and edge servers. This tight integration allows the LAWN to respond seamlessly to rapidly changing conditions, supporting applications such as swarm coordination, disaster response, urban surveillance, package delivery, and environmental monitoring.
@@ II-B How LAWN Differs from Aerial Communications?
Functional Scope Focused on communication only; separate control and sensing layers Integrated data, control, and sensing planes with a supporting I&C plane
@@
Computing Resources Centralized or onboard-only processing; limited scalability Hierarchical computing for real-time processing and decision-making
@@
In contrast, the LAWN introduces a paradigm shift by redefining the network around the terrestrial infrastructure, extending its functionality into the low-altitude space through a reconfigurable, intelligent, and mission-aware network “skyway”. Instead of simply using drones as communication tools, LAWN treats aerial and terrestrial nodes as jointly organized components of a cyber-physical system. The architecture tightly integrates three primary functional planes, i.e., data, control, and sensing, augmented by an auxiliary I&C plane. For example, traditional systems relying on fixed ground stations may struggle to adapt to such dynamic conditions. Dynamic reconfigurability of both aerial and terrestrial nodes sets LAWN apart from the configurations of traditional systems, ensuring robustness across diverse scenarios. A defining advantage of LAWN is its cooperative sensing capability, allowing aerial and ground nodes to share environmental data in real time. This shared perception supports safety-critical services such as multi-drone formation flying, dynamic obstacle avoidance, and urban air traffic management [^1] Traditional drone-assisted mesh networks, while capable of network-wide sharing, often lack sensing functionality supported by the sensing plane of the LAWN.. Additionally, LAWN’s auxiliary I&C plane provides distributed computational resources for real-time processing and decision-making. This includes on-board processing for immediate tasks, edge computing for latency-sensitive operations, and cloud-based infrastructure for intensive computations and long-term data analysis. Such a hierarchical framework is rarely embedded in traditional systems. In Table I, we summarize the comparison between the LAWN and the aerial communication system.